# QUOTA: Quantifying Objects with Text-to-Image Models for Any Domain

### v1

@@ 4.2 Dual-loop meta-optimization
QUOTA leverages a dual-loop meta-learning approach to achieve domain-invariant object quantification across diverse domains. This process involves an inner loop, which adapts the prompt parameters across multiple meta-train domains, and an outer loop, which refines these parameters on a single meta-test domain to encourage domain generalization. Each loop’s loss function incorporates two components: a quantification loss for object accuracy and a CLIP matching penalty to maintain semantic consistency. See Figure 2 for an overview.
@@
Inner loop optimization. In the inner loop, given a set of meta-train domains $\mathcal{D}_{\text{train}}$, we optimize the prompt parameters $\theta$ across multiple domains to ensure the adaptability of the model. For each domain $d\in\mathcal{D}_{\text{train}}$, the composite loss function includes:
@@
Quantification loss. Our inner loop optimization process utilizes a differentiable counting function, Count, which estimates the quantity of objects for a specified class in an image. The quantification loss for a target class $c$ is defined as the difference between the estimated and desired object counts, given by:
$\mathcal{L}_{\text{counting}}(x_{d},c,N)=\left\|\texttt{Count}(x_{d},c)-N\right\|,$ (5)
@@
where $\texttt{Count}(x_{d},c)$ provides an estimate of the object count for class $c$ within the generated image $x_{d}$. This loss is calculated using a detection-based dynamic scale, leveraging YOLOv9 [62] to estimate the count of objects of class $c$ in the generated image $x_{d}$. The estimated count is given by:
$\text{Count}(x_{d},c)=\sum_{i=1}^{h\cdot w}\hat{\lambda}_{\text{scale},x_{d},\mathcal{D}}\cdot\Phi_{i}(x_{d},c),$ (6)
@@
where $\hat{\lambda}_{\text{scale},x_{d}}$ is a dynamically adjusted scaling factor for differentiation continuity, and $\Phi_{i}(x_{d},c)$ represents the detection score for class $c$ at each pixel. This loss penalizes the difference between the target count $N$ and the detected count, ensuring accurate object quantification.
@@
CLIP matching penalty. To ensure the image semantics remain intact, we add a CLIP-based penalty term:
$\mathcal{L}_{\text{semantic}}(x_{d},c;\theta)=\text{CLIP}(x_{d},\text{``A }s_{d}\text{ of }c\text{''}),$ ” ) , (7)
@@
where $s_{d}$ is a domain descriptor specific to the domain $d$ (e.g., “photo,” “sketch,” “cartoon,” or “painting”), and CLIP provides a matching score between the generated image $x_{d}$ and the textual description “A [style] of [class],” reinforcing both semantic and stylistic alignment.
The inner loop loss $\mathcal{L}_{\text{inner}}$ is then computed by summing these losses across all meta-train domains:
$\mathcal{L}_{\text{inner}}(\theta)=\sum_{d\in\mathcal{D}_{\text{train}}}\mathcal{L}_{\text{counting}}(x_{d},c,N;\theta)+\lambda\mathcal{L}_{\text{semantic}}(x_{d},c;\theta),$ (8)
@@
Outer loop optimization. The outer loop further refines the prompt parameters by evaluating them on the meta-test domain $\mathcal{D}_{\text{test}}$ to promote generalization to unseen domains. After optimizing the prompt parameters $\theta$ in the inner loop across meta-train domains, the outer loop computes the loss on the meta-test domain, defined as:
$\mathcal{L}_{\text{outer}}(\phi)=\mathcal{L}_{\text{counting}}(x_{\text{test}},c,N;\theta)+\lambda\mathcal{L}_{\text{semantic}}(x_{\text{test}},c;\theta),$ (9)
where $x_{\text{test}}$ is the generated image using the updated prompt parameters $\theta$ from the inner loop on the meta-test domain $\mathcal{D}_{\text{test}}$, and $\phi$ represents the global meta-parameters. This outer optimization step updates $\phi$, ensuring that the learned prompt parameters are robust enough to generalize to new, unseen domains, achieving accurate object quantification while maintaining semantic fidelity.
Through this dual-loop meta-optimization framework, QUOTA adapts the prompt parameters to handle domain shifts effectively, enabling consistent object quantification across various domains without compromising image quality or semantic alignment.
@@ 4.3 Learnable quantification and domain tokens
To achieve adaptive control over both object quantification and domain in text-to-image generation, we introduce prompt learning with two learnable tokens, $e_{\text{count}}$ and $e_{\text{style}}$. These tokens enable the model to flexibly adjust to different domains while maintaining accurate object quantification. This approach enhances the model’s ability to generalize across diverse domains by encoding information specific to object quantification and stylistic features directly within the prompt. Notably, all other model parameters remain fixed, allowing the model to achieve domain adaptability without altering the underlying architecture.
The process starts with an initial prompt format: “A $s_{d}$ of $N$ $c$,” where $s_{d}$ represents the domain (e.g., “photo,” “sketch,” “cartoon”), $N$ specifies the target object count, and $c$ denotes the object class. To exercise control over the generated image, we concatenate two newly added pseudo token embeddings, $e_{\text{count}}$ and $e_{\text{style}}$, to the text embedding. These embeddings are randomly initialized and are iteratively updated during training to adaptively encode information about object quantification and domain.
Unlike optimizing existing tokens, which may corrupt tokens with multiple meanings or associations, the introduction of these pseudo tokens allows the model to learn domain-specific adjustments without impacting predefined tokens. Formally, we optimize the embeddings $e_{\text{count}}$ and $e_{\text{style}}$ for each image generated, where the prompt parameters $\theta$ in our meta-optimization framework refer to these learnable tokens. In the inner loop, $e_{\text{count}}$ and $e_{\text{style}}$ are trained across meta-train domains to capture domain-specific prompt characteristics, resulting in domain-specific embeddings. The outer loop then refines these embeddings by adjusting the meta-parameters to generalize across domains, using the meta-test domain to guide the final optimization. The outer loop’s updated meta-parameters represent our optimized, domain-generalizable prompt embeddings, ensuring consistency in object quantification and domain adaptation even for unseen domains.
@@ 4.4 Testing phase
During testing on a new domain, we simply incorporate the domain and quantification information from the test text prompt by concatenating it with our learned tokens, $e_{\text{style}}$ and $e_{\text{count}}$. For example, if the input text prompt is “a cartoon of 6 apples,” the prompt fed to SDXL [41] becomes “a cartoon $e_{\text{style}}$ of 6 $e_{\text{count}}$ apples.” This approach allows us to generate images accurately aligned with the desired domain and object count without needing to re-optimize $e_{\text{style}}$ and $e_{\text{count}}$, thus saving time and ensuring efficient adaptation to new domains.
@@
Table 1: Impact of dual-Loop meta-optimization on object quantification and semantic alignment across different domains on the FSC benchmark. Lower values for MAE and RMSE indicate better counting accuracy, while higher CLIP-S scores represent better alignment with prompt semantics. Meta-optimization shows improvements across all metrics, especially in challenging domains such as Sketch, demonstrating its effectiveness in enhancing both object quantification and domain generalization.
5 Experiments
5.1 Experimental setup
@@
w/ $e_{style}$ 14.41 19.26 73.3 11.31 15.28 73.7 12.73 15.07 72.9 9.91 13.25 72.7 12.28 16.15 73.2
w/ $e_{count}$ 14.12 19.60 73.9 10.20 14.44 74.1 10.66 13.17 73.7 8.75 11.89 73.3 10.94 14.76 73.8
Ours (w/ $e_{count}$ $e_{style}$) 10.09 13.98 74.3 8.04 11.04 75.9 9.48 11.92 74.8 9.15 11.69 74.1 9.19 12.16 74.8
@@
Table 2: Impact of learnable tokens on object quantification and semantic alignment across different domains on the FSC benchmark. Adding learnable tokens $e_{\text{count}}$ and $e_{\text{style}}$ improves YOLO-based MAE, RMSE, and CLIP-S scores, with the best performance achieved when both tokens are used together.
Figure 3: Comparison of object quantification results across domains with different optimization settings. This figure illustrates object quantification performance across various styles for three methods: SDXL [41], using only the inner loop in our framework, and our full QUOTA framework. The examples demonstrate how SDXL and the inner-loop-only setting struggle to achieve accurate quantification across domains, while meta-optimization achieves precise and consistent object quantification.
@@
Only test domain 5.68 8.63 75.9 4.61 7.12 76.8 5.31 7.86 75.9 5.94 8.53 75.3 5.38 8.04 76.0
Train domains & Test domain 6.43 9.05 74.7 7.61 10.79 76.1 8.43 10.96 75.1 8.75 11.27 74.7 7.81 10.52 75.2
Table 3: Performance comparison with various baseline models on object quantification and semantic alignment across different domains on the FSC benchmark. The methods and numbers shown in gray represent the upper bound, as these models have been trained using the test domain. Note that IoCo [66] is specifically trained on the Photo, which explains its high performance on Photo. In contrast, our QUOTA achieves strong results across all domains without having been explicitly trained on Photo, highlighting its domain-generalizable capabilities.
Figure 4: Generalization to unseen classes. Comparison of object quantification and style consistency between stable diffusion and our QUOTA, on both trained and unseen (reused) classes. While SDXL [41] struggles with unseen classes, QUOTA accurately controls object quantification and maintains semantic alignment across both trained classes (e.g., oranges, birds) and unseen classes (e.g., strawberries, crows). This demonstrates the robustness of QUOTA in adapting to previously unseen classes and domains.
Benchmark. We developed QUANT-Bench to assess object quantification performance across various classes, quantities, and visual domains. The classes in QUANT-Bench are based on the FSC dataset [47], a widely used dataset for object quantification that includes 19 categories such as common objects, vehicles, and animals. For each class, we generated 25 examples for quantities ranging from one to 25, yielding a total of 475 samples. To evaluate the model’s ability to generalize across different domains, we created images in four distinct domains: Photo, Painting, Cartoon, and Sketch. Both training and testing categories are drawn from the FSC dataset. To test each domain, we used a leave-one-out setting, where three domains are used as meta-train domains and the remaining domain as the meta-test domain. This setup allows us to measure how well the model generalizes object quantification to unseen domains by training with textual descriptions from the other three domains.
Metrics. To evaluate quantification accuracy, we use the MAE (Mean Absolute Error) and RMSE (Root Mean Square Error) metrics based on YOLOv9 [62]. These metrics are calculated by comparing the detected object counts to the target count in each image, with MAE capturing the average absolute deviation and RMSE highlighting larger errors through squared deviations. For semantic alignment assessment, we use a normalized CLIP matching score, defined as $\operatorname{CLIP-S}(p,x_{d})=w\cdot\max(\cos(p,x_{d}),0)$, where $p$ is the prompt and $x_{d}$ is the generated image [18]. Lastly, we calculate the average performance across different domains to assess each model’s consistency and adaptability to diverse contexts.
Implementation Details. We conducted training and evaluation on a single NVIDIA A6000 GPU with 48GB of memory. Training each token requires approximately two minutes, totaling around 30 iterations, using the SDXL framework [41]. For image quality, we find that a single denoising step is sufficient. The optimized quantification token can be reused without additional optimization time. We set $\lambda{=}5$, a learning rate of 0.01 for optimization, and the clipcount scaling hyperparameter $\lambda_{\text{scale}}$ to 60 for a static scale. We will make our code available.
Baselines. We establish a baseline using SDXL [41] backbone. We also compare with IoCo [66], which learns the quantification token only during training and considers images in the Photo domain exclusively.
5.2 Results
Benefit of meta-optimization. Table 1 demonstrates the impact of meta-optimization on object quantification accuracy and semantic alignment across different image styles. On average, meta-optimization achieves substantial improvements in both MAE and RMSE, highlighting enhanced object quantification accuracy and adaptability. The benefits are particularly pronounced in challenging domains, such as Cartoon, where meta-optimization leads to improved precision and consistency. Additionally, meta-optimization improves semantic alignment, as indicated by higher CLIP-S scores, reflecting closer alignment with prompt semantics. Figure 3 provides qualitative examples, comparing outputs from stable diffusion, an inner-loop-only optimization, and our full QUOTA framework. While inner-loop-only optimization offers some gains, it still struggles with consistent counts. In contrast, our complete meta-optimization approach, which includes both inner and outer loops, enables precise and domain-invariant object quantification across all styles, even in difficult cases like Sketch. These results confirm that meta-optimization significantly enhances the model’s ability to generalize quantification and semantic consistency across diverse domains.
Benefit of learnable tokens. Table 2 demonstrates the impact of adding learnable tokens $e_{\text{count}}$ and $e_{\text{style}}$ on object quantification accuracy and semantic alignment across different domains. The standard stable diffusion model, without any learnable tokens, shows limited adaptability, achieving lower accuracy across all domains. Introducing the $e_{\text{style}}$ token alone provides moderate improvements, indicating that adapting to domain style contributes positively to performance. Adding only the $e_{\text{count}}$ token leads to further enhancements, emphasizing the importance of object count adaptation. When both learnable tokens $e_{\text{count}}$ and $e_{\text{style}}$ are combined, as in QUOTA, the model achieves the best performance across all metrics and domains. This setup consistently enhances both counting accuracy and domain adaptability, highlighting the complementary effects of using both tokens. These results confirm that the joint use of $e_{\text{count}}$ and $e_{\text{style}}$ significantly strengthens the model’s ability to generalize object quantification across diverse domains while maintaining robust semantic alignment.
Figure 5: Qualitative comparison of generated images across different domains and classes. This comparison highlights the superior quantification accuracy and domain consistency achieved by QUOTA across diverse domains. The images outlined in red represent cases where the generated images have incorrect object quantification. The box in the Sketch domain showcases a challenging failure case for QUOTA, emphasizing the difficulty of accurately quantifying objects in highly stylized and intricate settings.
Generalization on unseen classes. In this experiment, we demonstrate the ability of our model, trained on 19 predefined classes, to generalize effectively to unseen classes using the optimized prompts generated by QUOTA. As illustrated in Figure 4, we tested the model on new classes not encountered during training, such as strawberries and crows. This generalization capability arises from our dual-loop training approach, where both the inner and outer loops are exposed to different domains and classes. By alternating between meta-train and meta-test domains with varying classes in each loop, our framework mimics the testing scenario of unseen class predictions, preparing the model for diverse and previously unseen content. Our approach successfully adapts to both the unseen classes and their respective domains, accurately controlling the object quantification and preserving semantic alignment. These results highlight QUOTA’s robustness in generalizing object quantification and domain consistency to new classes, demonstrating its potential for broad applicability in diverse scenarios.
@@ 5.2 Results
Comparison with different baselines. To provide a comprehensive understanding of QUOTA ’s performance, we include a qualitative comparison in Figure 5 alongside the quantitative results in Table 3. Figure 5 illustrates generated images across various domains and classes, with QUOTA demonstrating superior quantification accuracy and domain consistency. In contrast, SDXL [41] and IoCo [66] struggle with incorrect object quantification across different domains, as indicated by the red outlines on images with quantification errors. QUOTA achieves consistent object counts close to the target quantities in most cases, while the baseline models frequently misestimate counts, especially in complex domains. Notably, IoCo’s reliance on training within the photo domain explains its higher performance in this domain but highlights its limitations in generalizing to unseen styles. The highlighted box in the Sketch domain shows a failure case for QUOTA, emphasizing the challenges in accurately quantifying objects within highly stylized and intricate settings. Together, these visual and numerical comparisons underscore QUOTA ’s capability to set a new benchmark for object quantification accuracy and domain adaptability in text-to-image generation.
Limitations and Future Work. While our approach demonstrates robust performance across various domains, it has limitations in particularly challenging domains, where improvements are less pronounced. This may be due to the inherent difficulty of accurately quantifying objects in abstract or highly stylized representations, although our method still outperforms previous approaches in these cases. Additionally, the use of a dual-loop meta-learning strategy introduces a higher computational cost during training, as it requires more iterations for optimizing domain-invariant prompts. However, at test time, our approach remains faster than traditional adaptation-based methods, such as IoCo [66], which require additional adaptation steps.
For future work, we plan to explore alternative object detectors to improve object quantification accuracy in challenging domains. Additionally, our method could be extended to other text-to-image generation models and applied to different tasks, such as fine-grained attribute control and scene layout generation, broadening its applicability within generative modeling.

### v2

@@ 4.2 Dual-loop meta-optimization
QUOTA employs a dual-loop meta-optimization framework to achieve domain-invariant object quantification across diverse domains. This framework consists of an inner loop, which adapts the prompt parameters across multiple meta-train domains, and an outer loop, which refines these parameters on meta-validation domains to improve generalization. By dynamically sampling meta-train and meta-validation domains during training, the model learns to adapt to domain shifts and ensure robust object quantification. See Figure 2 for an overview.
Meta-learning domain partitioning. We denote the meta-train and meta-validation domains as $S^{tr}$ and $S^{val}$, respectively. During training, we dynamically partition $\mathcal{S}$ into meta-train domains $\mathcal{S}^{\text{tr}}$ and meta-validation domains $\mathcal{S}^{\text{val}}$. The model first optimizes domain-specific prompt parameters $\theta_{s^{\text{tr}}}$} using $\mathcal{S}^{\text{tr}}$, refining object quantification within each domain. It is then validated on $\mathcal{S}^{\text{val}}$ to refine domain-invariant prompt parameters $\phi$, promoting generalization beyond seen distributions.
@@
Inner loop optimization. In the inner loop, given a set of meta-train domains $\mathcal{S}^{\text{tr}}$, we optimize the prompt parameters $\theta$ across multiple domains to ensure the adaptability of the model. For each domain $s^{\text{tr}}\in\mathcal{S}^{\text{tr}}$, the composite loss function includes: quantification loss $\mathcal{L}_{\text{counting}}$ and CLIP matching penalty $\mathcal{L}_{\text{semantic}}$.
@@
Quantification loss. Our inner loop optimization uses a differentiable counting function, Count, which estimates the quantity of objects for a specified class in an image. The quantification loss for a target class $c$ is defined as the difference between the estimated and desired object counts, given by:
$\mathcal{L}_{\text{counting}}(x_{s}^{\text{tr}},c,N)=\left\|g_{\text{cnt}}(x_{s}^{\text{tr}},c)-N\right\|,$ (7)
@@
where $g_{\text{cnt}}(x_{s}^{\text{tr}},c)$ provides an estimate of the object count for class $c$ within the generated image $x_{s}^{\text{tr}}$. This loss is calculated using a detection-based dynamic scale, leveraging YOLOv9 [33] to estimate the count of objects of class $c$ in the generated image $x_{s}^{\text{tr}}$. The estimated count is given by:
$\text{Count}(x_{s}^{\text{tr}},c)=\sum_{i=1}^{h\cdot w}\hat{\lambda}_{\text{scale},x_{s}^{\text{tr}},\mathcal{D}}\cdot\Phi_{i}(x_{s}^{\text{tr}},c),$ (8)
@@
where $\hat{\lambda}_{\text{scale},x_{s}^{\text{tr}},\mathcal{D}}$ is a dynamically adjusted scaling factor for differentiation continuity, and $\Phi_{i}(x_{s}^{\text{tr}},c)$ represents the detection score for class $c$ at the $i$-th bounding box [35]. This loss penalizes the difference between the target count $N$ and the detected count, ensuring accurate object quantification.
@@
CLIP matching penalty. To preserve image semantics, we add a CLIP-based penalty term:
$\mathcal{L}_{\text{semantic}}(x_{s}^{\text{tr}},c,N)=\text{CLIP}(x_{s}^{\text{tr}},\text{``A }s_{d}\text{ of }N\text{ }c\text{''}),$ (9)
@@
where $s_{d}$ is a domain descriptor specific to the domain $s$ (e.g., “photo,” “sketch,” “cartoon,” or “painting”), and CLIP provides a matching score between the generated image $x_{s}^{\text{tr}}$ and the textual description “A [style] of N [class],” reinforcing both semantic and stylistic alignment. The inner loop loss $\mathcal{L}_{\text{inner}}$ is then computed by summing these losses across all meta-train domains:
$\mathcal{L}_{\text{inner}}({\theta_{s^{\text{tr}}}})=\sum_{s^{\text{tr}}\in\mathcal{S}^{\text{tr}}}\mathcal{L}_{\text{counting}}(x_{s}^{\text{tr}},c,N;{\theta_{s^{\text{tr}}}})$ (10)
$+\lambda\mathcal{L}_{\text{semantic}}(x_{s}^{\text{tr}},c,N;{\theta_{s^{\text{tr}}}}),$ (11)
@@
Outer loop optimization. The outer loop further refines the prompt parameters by evaluating them on the meta-validation domain $\mathcal{S}^{\text{val}}$ to promote generalization to unseen domains. After optimizing the prompt parameters $\theta_{s^{\text{val}}}^{*}$} in the inner loop across meta-train domains, the outer loop computes the loss on the meta-validation domain, defined as:
$\mathcal{L}_{\text{outer}}(\phi)=\mathcal{L}_{\text{counting}}(x^{\text{val}}_{s},c,N;{\theta_{s^{\text{val}}}^{*}})$ (12)
$+\lambda\mathcal{L}_{\text{semantic}}(x^{\text{val}}_{s},c,N;{\theta_{s^{\text{val}}}^{*}}),$ (13)
where $x^{\text{val}}_{s}$ is the generated image using the updated prompt parameters $\theta_{s^{\text{val}}}^{*}$} from the inner loop, and $\phi$ denotes the shared meta-parameters, i.e., the global initialization of $e_{\text{count}}$ and $e_{\text{style}}$. The inner loop adapts these tokens to domain-specific requirements, while the outer loop updates their shared initialization to improve generalization across domains. Unlike the inner loop, where the text prompt follows the structure “A $s_{d}^{\text{tr}}$ of N cc”, the outer loop utilizes a different domain descriptor derived from $\mathcal{S}^{\text{val}}$, modifying the prompt format to “A $s_{d}^{\text{val}}$ of N cc”. This ensures that the model is exposed to different domain styles between training and validation, reinforcing robustness to domain shifts. By training on $\mathcal{S}^{\text{tr}}$ and validating on $\mathcal{S}^{\text{val}}$, the outer optimization step updates $\phi$, ensuring that the learned prompt parameters generalize effectively to new, unseen domains. This process achieves accurate object quantification while maintaining semantic fidelity. Through this dual-loop meta-optimization framework, QUOTA adapts the prompt parameters to handle domain shifts effectively, enabling consistent object quantification across various domains without compromising image quality or semantic alignment.
@@ 4.3 Learnable quantification and domain tokens
To achieve adaptive control over both object quantification and the domain in text-to-image generation, we introduce prompt learning with two learnable tokens, $e_{\text{count}}$ and $e_{\text{style}}$. These tokens enhance the model’s ability to generalize across diverse domains by encoding information specific to object quantification and stylistic features directly within the prompt. Notably, all other model parameters remain fixed, allowing the model to achieve domain adaptability without altering the underlying architecture. To exercise control over the generated image, we concatenate two newly added pseudo token embeddings, $e_{\text{count}}$ and $e_{\text{style}}$, to the text embedding. These embeddings are randomly initialized and are iteratively updated during training to adaptively encode information about object quantification and domain. Unlike optimizing existing tokens, which may corrupt tokens with multiple meanings or associations, the introduction of these pseudo tokens allows the model to learn domain-specific adjustments without impacting predefined tokens. Formally, we optimize the embeddings $e_{\text{count}}$ and $e_{\text{style}}$ for each image generated, where the prompt parameters $\theta$ in our meta-optimization framework refer to these learnable tokens. In the inner loop, $e_{\text{count}}$ and $e_{\text{style}}$ are trained across meta-train domains to capture domain-specific prompt characteristics, resulting in domain-specific embeddings. The outer loop then refines these embeddings by adjusting the meta-parameters to generalize across domains, using the meta-validation domain to guide the final optimization. The outer loop’s updated meta-parameters represent our optimized, domain-generalizable prompt embeddings, ensuring consistency in object quantification and domain adaptation even for unseen domains.
@@ 4.4 Testing phase
During testing on a new domain $\mathcal{T}$, we simply incorporate the domain and quantification information from the test text prompt by concatenating it with our learned tokens, $e_{\text{style}}$ and $e_{\text{count}}$. For example, if the input text prompt is “a cartoon of 6 apples,” the prompt fed to SDXL [23] becomes “a cartoon $e_{\text{style}}$ of 6 $e_{\text{count}}$ apples.” This approach allows us to generate images accurately aligned with the desired domain and object count without needing to re-optimize $e_{\text{style}}$ and $e_{\text{count}}$, thus saving time and ensuring efficient adaptation to new domains.
@@
Table 1: Benefit of meta-optimization on object quantification and semantic alignment across different domains on QUANT-Bench. Meta-optimization shows improvements across all metrics, especially in challenging domains such as Sketch, demonstrating its effectiveness in enhancing both object quantification and domain generalization.
@@
w/ $e_{\text{style}}$ 14.41 19.26 73.3 11.31 15.28 73.7 12.73 15.07 72.9 9.91 13.25 72.7 12.28 16.15 73.2
w/ $e_{\text{count}}$ 14.12 19.60 73.9 10.20 14.44 74.1 10.66 13.17 73.7 8.75 11.89 73.3 10.94 14.76 73.8
Ours (w/ $e_{\text{count}}$ & $e_{\text{style}}$) 10.09 13.98 74.3 8.04 11.04 75.9 9.48 11.92 74.8 9.15 11.69 74.1 9.19 12.16 74.8
@@
Table 2: Benefit of learnable tokens on object quantification and semantic alignment across domains on QUANT-Bench. Adding learnable tokens $e_{\text{count}}$ and $e_{\text{style}}$ shows improvements across all metrics, with the best performance achieved when both tokens are used together.
5 Experiments
5.1 Experimental setup
QUANT-Bench. We developed QUANT-Bench to assess object quantification performance across various classes, quantities, and visual domains. The classes in QUANT-Bench are based on the FSC dataset by Ranjan et al. [25], a widely used dataset for object counting that includes 19 classes such as common objects, vehicles, and animals. For each class, we generated 25 examples for quantities ranging from one to 25, yielding a total of 475 samples. To evaluate the model’s ability to generalize across different domains, we created images in four distinct domains: Photo, Painting, Cartoon, and Sketch. Both training and testing classes are drawn from the FSC dataset. To test each domain, we used a leave-one-domain-out setting, where two domains are used as meta-train domains for the inner loop, one domain is used as the meta-validation domain for the outer loop, and the remaining domain is held out as the unseen test domain. This setup allows us to measure how well the model generalizes object quantification to unseen domains by training with textual descriptions from the other three domains.
Metrics. To evaluate quantification accuracy, we use the Mean Absolute Error (MAE) and the Root Mean Squared Error (RMSE) based on YOLOv9 [33]. These metrics are calculated by comparing the detected object counts to the target count in each image, with MAE capturing the average absolute deviation and RMSE highlighting larger errors through squared deviations. For semantic alignment assessment, we use the CLIP-S from Eq. (9), following [10]. Conventional metrics such as FID [11] and Inception Score [28] are not appropriate for our QUANT-Bench. FID requires a real reference distribution, which is not available in our synthetic prompt space, and the Inception Score emphasizes sample diversity rather than alignment with the input prompt.
Implementation details. We conducted training and evaluation on a single NVIDIA A6000 GPU with 48GB of memory. Training each token requires approximately two minutes, totaling around 30 iterations, using the SDXL framework [23]. For image quality, we find a single denoising step suffices. The optimized quantification token can be reused without additional optimization time. We set $\lambda{=}5$, a learning rate of 0.01 for optimization, and the scaling hyperparameter $\lambda_{\text{scale}}$ to 60 for a static scale. Code will be released.
5.2 Results
Benefit of meta-optimization. Table 1 demonstrates the impact of meta-optimization on object quantification accuracy and semantic alignment across different image styles. On average, meta-optimization achieves substantial improvements in both MAE and RMSE, highlighting enhanced object quantification accuracy and adaptability. The benefits are particularly pronounced in challenging domains, such as Cartoon, where meta-optimization leads to improved precision and consistency. Additionally, meta-optimization improves semantic alignment, as indicated by higher CLIP-S scores, reflecting closer alignment with prompt semantics. Figure 3 provides qualitative examples, comparing outputs from stable diffusion, an inner-loop-only optimization, and our full QUOTA framework. While inner-loop-only optimization offers some gains, it still struggles with consistent counts. In contrast, our complete meta-optimization approach, which includes both inner and outer loops, enables precise and domain-invariant object quantification across all styles, even in difficult cases like Sketch. These results confirm that meta-optimization considerably enhances the model’s ability to generalize quantification and semantic consistency across diverse domains.
Benefit of learnable tokens. Table 2 demonstrates the impact of adding learnable tokens $e_{\text{count}}$ and $e_{\text{style}}$ on object quantification accuracy and semantic alignment across different domains. The standard SDXL, without any learnable tokens, shows limited adaptability, achieving lower accuracy across all domains. Introducing the $e_{\text{style}}$ token alone provides moderate improvements, indicating that adapting to domain style contributes positively to performance. Adding only the $e_{\text{count}}$ token leads to further enhancements, emphasizing the importance of object count adaptation. When both learnable tokens $e_{\text{count}}$ and $e_{\text{style}}$ are combined, as in QUOTA, the model achieves the best performance across all metrics and domains. These results confirm that the joint use of $e_{\text{count}}$ and $e_{\text{style}}$ strengthens the model’s ability to generalize object quantification across diverse domains while maintaining robust semantic alignment.
Figure 3: Benefit of meta-optimization. We compare SDXL [23], our method with only the inner loop, and full QUOTA. Without meta-optimization, models struggle with accurate quantification. QUOTA achieves consistent counting across domains.
Figure 4: Generalization to unseen classes. Comparison of SDXL [23] and QUOTA on both seen and unseen classes. QUOTA maintains accurate quantification and style consistency, while SDXL struggles with unseen classes, highlighting our method’s robustness.
@@ 5.2 Results
Target Domain 5.68 8.63 75.9 4.61 7.12 76.8 5.31 7.86 75.9 5.94 8.53 75.3 5.38 8.04 76.0
Train & Target Domain 6.43 9.05 74.7 7.61 10.79 76.1 8.43 10.96 75.1 8.75 11.27 74.7 7.81 10.52 75.2
Table 3: Comparison with others on object quantification and semantic alignment across domains in QUANT-Bench. Gray numbers denote target-domain upper bounds. While IoCo [35] is trained on Photo, QUOTA performs well across all domains without such supervision, highlighting its generalizability.
Photo Painting Cartoon Sketch Average
SDXL [23] 14.07 19.80 72.8 20.64 27.05 73.0 11.71 15.40 72.1 18.39 24.03 71.1 16.20 21.57 72.3
IoCo [35] 13.42 18.97 73.3 13.16 19.01 73.4 11.29 14.73 72.7 14.41 19.71 71.60 13.07 18.11 72.8
QUOTA 13.72 19.20 73.1 10.99 15.24 74.5 9.94 12.22 73.2 12.05 14.68 72.3 11.68 15.33 73.3
Table 4: Generalization to unseen classes. Comparison of SDXL [23], IoCo [35], and QUOTA on unseen classes across four target domains. QUOTA achieves consistently lower MAE and RMSE together with higher CLIP-S scores, demonstrating stronger quantification accuracy, semantic alignment, and robustness to domain shifts.
Figure 5: Comparison with others by their generated images across different domains and object classes. This figure highlights the quantification accuracy and domain consistency of QUOTA. In the Painting domain (apples), QUOTA generates seven apples, but one appears significantly smaller, making it difficult for YOLOv9 to detect. Similarly, in the Sketch domain, the highly stylized and intricate setting challenges accurate object quantification, illustrating the limitations of detection-based counting in complex visual styles.
Generalization to unseen classes. We further evaluate the ability of QUOTA, trained on 19 predefined classes, to generalize to classes not seen during training. Figure 4 provides qualitative examples, where we test on new categories such as strawberries and crows. While SDXL [23] struggles to maintain correct counts and style consistency, QUOTA successfully adapts to these unseen categories and domains, producing accurate quantification and faithful semantic alignment. Table 4 complements this analysis with quantitative results across four target domains. The unseen classes are grouped by semantic similarity to training classes (e.g., apples generalize to tomatoes, oranges, and strawberries; birds to crows, pigeons, and seagulls; sheep to zebras, horses, and cows). We report MAE, RMSE, and CLIP-S scores averaged over nine unseen classes with quantities from 1 to 25. Across all domains (Photo, Painting, Cartoon, and Sketch), QUOTA consistently outperforms SDXL [23] and IoCo [35], achieving substantially lower MAE and RMSE together with higher CLIP-S scores. These qualitative and quantitative results together highlight QUOTA’s robustness in extending object quantification to unseen classes while preserving semantic consistency, demonstrating its broad applicability across diverse domains and distributions.
Methods MAE $\downarrow$ RMSE $\downarrow$ CLIP-S $\uparrow$
Simple 9.19 12.16 74.8
Complex 7.23 9.01 73.1
Simple $\rightarrow$ complex 8.25 10.05 73.8
Table 5: Effect of prompt complexity. Simple uses default prompts, while Complex includes GPT-4o-generated descriptions and trains the count token. Applying tokens from Simple to Complex (Simple $\rightarrow$ Complex) shows our method’s robustness to prompt variation.
Effect of prompt complexity. To evaluate the impact of prompt complexity on object quantification, we compare three settings: Simple, Complex, and Simple $\rightarrow$ Complex, as shown in Table 5. The Simple setting uses default prompts, while Complex incorporates GPT-4o-generated [21] detailed descriptions and trains the count token as a learnable token, leading to improved performance. To assess the adaptability of our method, we apply the style and count tokens trained on simple prompts directly to complex prompts, which results in performance of the "Simple to Complex" setting. This demonstrates our method effectively generalizes across prompt complexities, maintaining robust object quantification without requiring additional retraining.
@@
Comparison with others. To provide a comprehensive understanding of QUOTA’s performance, we include a qualitative comparison in Figure 5 alongside the quantitative results in Table 3. Figure 5 illustrates generated images across various domains and classes, with QUOTA demonstrating superior quantification accuracy and domain consistency. In contrast, both SDXL [23] and IoCo [35] struggle with incorrect object quantification across different domains, as indicated by the red outlines on images with quantification errors. QUOTA achieves consistent object counts matching the target quantities in most cases, while the baseline models frequently misestimate counts, especially in complex domains. Notably, IoCo’s reliance on training within the photo domain explains its higher performance in this domain but highlights its limitations in generalizing to unseen styles. Together, these visual and numerical comparisons underscore QUOTA’s capability to set a new benchmark for object quantification accuracy and domain adaptability in text-to-image generation.